OpenStudy (anonymous):

PLEASE HELP ME! The sum of three numbers is 27. The second number is two less than twice the first number, and the third number is five greater than three times the first

OpenStudy (anonymous):

Pls help me.

OpenStudy (cj49):

let the 3 no be x,y and z x+y+z=27

OpenStudy (anonymous):

okay. and then?

OpenStudy (cj49):

the 2nd no is 2 less then twice the 1st no y=2x-2

OpenStudy (cj49):

the 3rd no is 5 greater than 3 times the 1st no z=3x+5

OpenStudy (cj49):

x+(2x-2)+3x+5=27

OpenStudy (ahsome):

The sum of three numbers is 27. The second number is two less than twice the first number, and the third number is five greater than three times the first Let the first number be \(a\) Let the second number be \(b\) Let the third number be \(c\) \(a=a\) \(b=2a-2\) \(c=3a+5\) That is from your info \[a+b+c=27\] Sub in the values above \[a+b+c=27\] \[a+2a-2+3a+5=27\] Find \(a\) \[5a-2+5=27\] \[5a+3=27\] \[5a=24\] \[a=4.8\] Now sub the \(a\) value to get \(b\) and \(c\) \[b=2a-2\]\[b=2*4.8-2\]\[b=7.6\] \[c=3a+5\]\[c=3*4.8+5\]\[c=19.4\]

OpenStudy (ahsome):

Phew, that was a long one ;) Do you understand, @Socallmesheckyy?

OpenStudy (cj49):

x+2x-2+3x+5=27 6x+3=27 6x=24 x=4 x=4 y=6 z=17

OpenStudy (ahsome):

Whoops, I put 5a, not 6a. Just put 6 instead. Sorry :(

OpenStudy (anonymous):

THANK YOU @Ahsome , Ill try to understand that one. :) thanks to you too, @cj49

OpenStudy (ahsome):

I can rewrite if you want.

OpenStudy (cj49):

np

OpenStudy (ahsome):

No problem, @Socallmesheckyy